# A Roadmap for Simulating Chemical Dynamics on a Parametrically Driven Bosonic Quantum Device

### v1

@@ Abstract
In transition-state theory (TST),1, 2, 3 the reaction rate constant is given by the Eyring-Polanyi equation: $k_{TST}=\frac{k_{B}T}{h}e^{-\Delta G^{\ddagger}/k_{B}T}$. Here, $\Delta G^{\ddagger}$ is the Gibbs free energy of activation (the barrier height as measured from the minimum of the reactant well to the transition state configuration that corresponds to the barrier top), $k_{B}$ is the Boltzmann constant, $T$ is the absolute temperature, and $h$ is the Plank constant. The Gibbs free energy of activation can be divided into enthalpy and entropy contributions: $\Delta G^{\ddagger}=\Delta H^{\ddagger}-T\Delta S^{\ddagger}$, where $\Delta H^{\ddagger}$ is the activation enthalpy and $\Delta S^{\ddagger}$ is the activation entropy. Identifying $\Delta H^{\ddagger}$ as the activation energy ($E_{a}$), the TST rate constant can be expressed by Arrhenius’ law, $k_{TST}=A~{}e^{-E_{a}/k_{B}T}$, with $A=\frac{k_{B}T}{h}e^{\Delta S^{\ddagger}/k_{B}}$ the rate constant for the barrierless case (i.e., when $\Delta H^{\ddagger}=E_{a}=0$).
@@
In this paper, we propose a novel strategy for simulating the dynamics of elementary reactions using a quantum platform. We explore the superconducting circuit quantum electrodynamics (cQED) Kerr-cat device. 28, 29, 30, 31, 32, 33, 34, 25 We begin by introducing the Kerr-cat device and highlighting the features that make it ideally suited for analog simulations of chemical dynamics. Next, we consider examples of molecular systems to demonstrate its capabilities as applied to simulating proton transfer dynamics.
@@
Here, $\hat{a}^{\dagger}$ and $\hat{a}$ are the ladder operators for the device bosonic mode, satisfying the usual bosonic commutator relation $[\hat{a},\hat{a}^{\dagger}]=1$. The Hamiltonian $\hat{H}_{KC}$ includes adjustable parameters, namely the Kerr non-linearity, $K$, the detuning parameter, $\Delta$, and the drive coefficients, $\epsilon_{1}$ and $\epsilon_{2}$.
Figure 1: Schematic of Bosonic cQED Device for Quantum Dynamics Simulations in Hydrogen-Bonded Complexes (figure adapted from Ref. 25). The device is a half-aluminum, half-copper cavity resonator containing two sapphire chips with a SNAIL-transmon, readout resonator, and Purcell filter (top left). A strong microwave drive at twice the resonance frequency converts the SNAIL-transmon Hamiltonian into the bi-stable Kerr-cat parametric oscillator Hamiltonian for analog simulations of quantum dynamics in molecular systems, such as the adenine-thymine dimer (top right). Parameters $\Delta,K,\epsilon_{1},\epsilon_{2}$ (middle left) are adjusted to map the effective Hamiltonian $\hat{H}_{KC}$ to the double-well Hamiltonian modeling the molecular system (middle right). The bottom left shows the Wigner transform phase space representation of the device quantum states, which are analogous to the states in molecular the double-well potential in the bottom right.
Following Venkatraman et al.29 we assume that the dynamics of the device can be described by the following Lindblad equation:
@@
where $\kappa$ sets the photon loss rate corresponding to an effective dissipation rate due to coupling of the system with a surrounding environment and $n_{\text{th}}$ defines the effective temperature. Like $\{K,\Delta,\epsilon_{1},\epsilon_{2}\}$, the parameters $\kappa$ and $n_{\text{th}}$ are experimentally tunable. 36
@@
which is commonly used to simulate hydrogen bonded complexes. 18 The position and momentum operators of the reaction coordinate, $\hat{x}$ and $\hat{p}$, satisfy the commutation relation $[\hat{x},\hat{p}]=i\hbar$. In the equation above, $m$ represents the effective mass associated with motion along the reaction coordinate. The parameters $\{k_{1},k_{2},k_{4}\}$ are positive and real, typically determined by fitting the ab initio potential energy surface. This fitting process ensures the accurate representation of the barrier height, the curvature of the surface at the reactant and product wells, and the relative stability of reactants and products as parametrized by $k_{1}$. Table 1 lists the parameters $\{k_{1},k_{2},k_{4}\}$ used in this study for simulations of proton transfer dynamics (with $m=1836$ amu) in the four model systems illustrated in Fig. 2.
@@
Adenine-Thymine (DNA) 37 $1.4\times 10^{-3}$ $1.08\times 10^{-2}$ $5.2\times 10^{-3}$
Guanine-Cytosine (DNA) 38 $7.7\times 10^{-4}$ $6.9\times 10^{-3}$ $4.5\times 10^{-3}$
Malonaldehyde (cis-trans) 39, 40 $9.4\times 10^{-5}$ $3.0\times 10^{-3}$ $2.9\times 10^{-3}$
Malonaldehyde (cis-cis) 39, 40 $7.1\times 10^{-4}$ $4.0\times 10^{-3}$ 0
@@
Figure 2: Hydrogen bonded complexes studied to analyze intramolecular proton transfer in cis-cis and cis-trans malonaldehyde (top), intramolecular proton transfer in adenine-thymine (amine, pyrimidine) (middle), and guanine-cytosine (amine, pyrimidine) base pairs.
@@
We now map the parameters $\{\Delta,K,\epsilon_{2},\epsilon_{1}\}$ onto $\{m,k_{1},k_{2},k_{4}\}$ by ensuring consistency between the $\hat{p}^{2}$, $\hat{x}^{4}$, $\hat{x}^{2}$ and $\hat{x}$ terms in Eqs. (3) and (A Roadmap for Simulating Chemical Dynamics on a Parametrically Driven Bosonic Quantum Device). This leads to the following mapping relations:
@@
Clearly, according to Eqs. (6)-(9), the values of $\{K,\Delta,\epsilon_{1},\epsilon_{2}\}$ depend on the value of $c$. We then utilize the flexibility in choosing the value of $c$ to minimize the effect of the additional terms $\hat{p}^{4}$, $\hat{x}^{2}\hat{p}^{2}$ and $\hat{x}^{2}\hat{p}^{2}$ in Eq. (3). As demonstrated in the SI, this requires choosing a value of $c$ small enough to satisfy the following inequality:
@@
Using the cQED device to encode $\hat{H}_{\mathrm{DW}}$ requires a finite, non-zero value for the Kerr non-linearity $K$, so the units of $\Delta$, $\epsilon_{2}$ and $\epsilon_{1}$ are expressed in terms of $K$. For vanishingly small values of $c$, the ratios $\frac{\Delta}{K}$,$\frac{\epsilon_{2}}{K}$, and $\frac{\epsilon_{1}}{K}$ become very large (panels II of Fig. 4), so they are experimentally unfeasible with current cQED platforms. 41 However, there exists a range of values of $c$ for which the parameters are experimentally accessible, while maintaining a useful degree of accuracy for the energies and stationary states. In particular, the Hamiltonian for the cis-cis malonaldehyde is both experimentally accessible and meets the chemical accuracy criterion for stationary states and energies (see Fig. 4A). Chemical systems with asymmetric free energy profiles pose greater challenges due to the higher number of eigenstates required for addressing kinetic questions. Expanding the range of experimental parameters could enable an accurate simulation of these more challenging double-well problems, although it may also risk a breakdown of the effective Hamiltonian approximation.41
@@
Consistency between the energy levels of the $\hat{H}_{KC}$ and $\hat{H}_{DW}$ Hamiltonians is necessary, but not sufficient to ensure that simulations on the Kerr-cat device accurately capture the chemical dynamics. This is because the reaction rate is also influenced by the coupling of the reaction coordinate to the thermal bath of nonreactive DOF. Consequently, the actual reaction rate constant depends not only on the eigenvalues and eigenfunctions of the Hamiltonian but also on the temperature and dissipation rate of the bath. Therefore, being able to simulate chemical dynamics on the Kerr-cat device requires that the dynamics of the device be dissipative, which is indeed the case. The Kerr-cat device is described by the Lindblad equation in Eq. (2) with adjustable photon loss rate constant $\kappa$ and temperature set by $n_{th}$.
@@
The initial state for dynamics propagation is localized in the reactant well of the above-mentioned reactions. To generate this state, we diagonalize the system Hamiltonian and select the first eigenstate with more than 50% density in the reactant well. We then apply a sigmoidal filter function to remove the excess density outside of it (see SI for implementation details). This general localization method typically creates a localized Gaussian state and avoids the need to identify the critical points of each potential energy surface. By integrating Eq. (A Roadmap for Simulating Chemical Dynamics on a Parametrically Driven Bosonic Quantum Device) with an initial state $\hat{\rho}(0)$ localized in the reactant well, we obtain the time-evolved state $\hat{\rho}(t)$, which can be used to compute the product population at time $t$, as follows:
@@
The reactant-to-product reaction rate constant, $k=1/T_{X}$, is obtained by fitting $P_{P}(t)$ to an exponential decay, as described below. The top subpanels of Fig. 5 show the time evolution of $P_{P}(t)$ for the four proton transfer reactions – cis malonaldehyde, cis-trans malonaldehyde, adenine-thymine, and guanine-cytosine – obtained by solving Eq. (A Roadmap for Simulating Chemical Dynamics on a Parametrically Driven Bosonic Quantum Device) for different values of $c$ (solid lines). These top subpanels also include results where the chemical double-well Hamiltonian in Eq. (A Roadmap for Simulating Chemical Dynamics on a Parametrically Driven Bosonic Quantum Device), $\hat{H}_{DW}$, is replaced by the corresponding device Hamiltonian $-\hat{H}_{KC}$ with additional terms $\hat{p}^{4}$, $\hat{x}^{2}\hat{p}^{2}$ and $\hat{p}^{2}\hat{x}^{2}$ (diamonds). The bottom subpanels of Fig. 5 show the dependence of the inverse reaction rate constant, $T_{X}$, on $c$ for both $\hat{H}_{DW}$ (solid line) and $-\hat{H}_{KC}$ (diamonds).
A close examination of the results in Fig. 5 reveals that the dynamics of both the chemical system and the Kerr-cat device align well with the rate kinetics described by a rate constant. Additionally, the reaction rate constant shows significantly less sensitivity to the value of $c$ compared to the energy levels and eigenfunctions of $\hat{H}_{DW}$ and $\hat{H}_{KC}$. Specifically, the rate constant predicted by the device matches that predicted for the chemical system at values of $c$ as high as 0.4 $\mathrm{a_{0}}$, much higher than the value of $c=0.1\,\mathrm{a_{0}}$ required to match energy levels and eigenfunctions near the top of the barrier. This is likely due to the effect of dissipation, making the overall dynamics less sensitive to small differences in the potential energy surface.
@@
The results in Fig. 5 were obtained with $\kappa=0.1$ and $n_{\text{th}}=0.1$. However, the observed trends are insensitive to these specific values, as demonstrated by the rate constants obtained for various values of $\kappa$ and $n_{\text{th}}$ shown in Table 2 for $c=0.1\,\mathrm{a_{0}}$. As expected, $T_{X}$ increases (i.e., the reaction slows down) when $\kappa$ and $n_{\text{th}}$ decrease. Nonetheless, the actual values of $T_{X}$ for the chemical double-well and the Kerr-cat device are consistent, regardless of the $\kappa$ and $n_{\text{th}}$ values.
@@
Includes a detailed description of the methods and additional figures.
@@ 1 Code and Data Availability
The Python code for the Hamiltonian analysis and dynamical simulations can be found at https://github.com/dcabral00/cQED4ChemDyn. The data files used to generate the figures in the manuscript are also provided at https://github.com/dcabral00/cQED4ChemDyn and hosted at Zenodo.47
@@ 1.1 Mapping the Device Hamiltonian to Chemical Double-Well
We consider the effective Hamiltonian for currently experimentally realizable parametrically-driven Kerr-Cat cQED devices, which is given by:28, 29, 30, 31, 32, 48
@@
We also take advantage of the aforementioned flexibility in choosing the value of $c$ to minimize the effect of the spurious $\hat{p}^{4}$, $\hat{x}^{2}\hat{p}^{2}$ and $\hat{x}^{2}\hat{p}^{2}$ terms in Eq. (14). As we show below, doing so requires that we choose a value of $c$ small enough so that it satisfies the following inequality:
@@
The fact that the $\hat{p}^{4}$ term scales like $c^{4}$, while the $\hat{p}^{2}$ term scales like $c^{2}$, also implies that the $\hat{p}^{4}$ will becomes negligible for a sufficiently small value of $c$. In fact, the same inequality, Eq. (22), can be derived by nothing that $\frac{1}{2m}\hat{p}^{2}\gg\frac{E_{a}c^{8}}{\hbar^{4}x_{0}^{4}}\hat{p}^{4}$ is equivalent to $\frac{1}{2m}\gg\frac{E_{a}c^{8}}{\hbar^{4}x_{0}^{4}}\hat{p}^{2}$ and noting that the momentum is maximal when the particle is around the minima, where the potential energy can be approximated as being harmonic. Invoking the virial theorem for the harmonic oscillator, according to which the expectation values of the kinetic energy is equal to that of the potential energy, and noting that $E_{a}$ sets the potential energy scale for the chemical system, we can then estimate $\hat{p}^{2}$ by $2mE_{a}$ in the inequality $\frac{1}{2m}\gg\frac{E_{a}c^{8}}{\hbar^{4}x_{0}^{4}}\hat{p}^{2}$, which turns it into the inequality $\left(\frac{\hbar^{2}x_{0}^{2}}{2mE_{a}c^{4}}\right)^{2}\gg 1$. Thus, satisfying the inequality in Eq. (22), which is equivalent to $\frac{\hbar^{2}x_{0}^{2}}{2mE_{a}c^{4}}\gg 1$, also guarantees that the $\hat{p}^{4}$ term will becomes negligible compared the the $\hat{p}^{2}$ kinetic energy term.
Finally, the argument for why the same argument would also hold for an asymmetrical double-well has to do with the expectation that the length and energy scales of the chemical system are not going to be significantly affected by the addition of the asymmetry.

### v2

@@ Abstract
In transition-state theory (TST),1, 2, 3 the reaction rate constant is given by the Eyring-Polanyi equation: $k_{TST}=\frac{k_{B}T}{h}e^{-\Delta G^{\ddagger}/k_{B}T}$. Here, $\Delta G^{\ddagger}$ is the Gibbs free energy of activation (the barrier height as measured from the minimum of the reactant well to the transition state configuration that corresponds to the barrier top), $k_{B}$ is the Boltzmann constant, $T$ is the absolute temperature, and $h$ is the Planck constant. The Gibbs free energy of activation can be divided into enthalpy and entropy contributions: $\Delta G^{\ddagger}=\Delta H^{\ddagger}-T\Delta S^{\ddagger}$, where $\Delta H^{\ddagger}$ is the activation enthalpy and $\Delta S^{\ddagger}$ is the activation entropy. Identifying $\Delta H^{\ddagger}$ as the activation energy ($E_{a}$), the TST rate constant can be expressed by Arrhenius’ law, $k_{TST}=A~{}e^{-E_{a}/k_{B}T}$, with $A=\frac{k_{B}T}{h}e^{\Delta S^{\ddagger}/k_{B}}$ the rate constant for the barrierless case (i.e., when $\Delta H^{\ddagger}=E_{a}=0$).
@@
In this paper, we propose a novel strategy for simulating the dynamics of elementary reactions using a quantum platform. We explore the superconducting circuit quantum electrodynamics (cQED) Kerr-cat device. 28, 29, 30, 31, 32, 33, 34, 25 We begin by introducing the device and highlighting the features that make it ideally suited for analog simulations of chemical dynamics. Next, we consider examples of molecular systems to demonstrate the Kerr-cat device capabilities as applied to simulating proton transfer dynamics.
@@
Here, $\hat{a}^{\dagger}$ and $\hat{a}$ are the ladder operators for the device bosonic mode, satisfying the usual bosonic commutator relation $[\hat{a},\hat{a}^{\dagger}]=1$. The Hamiltonian $\hat{H}_{KC}$ includes adjustable parameters, namely the Kerr non-linearity, $K$, the detuning parameter, $\Delta$, and the drive coefficients, $\epsilon_{1}$ and $\epsilon_{2}$. These parameters can be experimentally altered by adjusting the magnetic flux and the amplitudes and frequencies of the microwave drives within the quantum device.36
Figure 1: Schematic of Bosonic cQED Device for Quantum Dynamics Simulations in Hydrogen-Bonded Complexes. Adapted from Ref. 25. Copyright (2024) American Chemical Society. The device is a half-aluminum, half-copper cavity resonator containing two sapphire chips with a SNAIL-transmon, readout resonator, and Purcell filter (top left).36 A strong microwave drive at twice the resonance frequency converts the SNAIL-transmon Hamiltonian into the bi-stable Kerr-cat parametric oscillator Hamiltonian for analog simulations of quantum dynamics in molecular systems, such as the adenine-thymine dimer (top right). Parameters $\Delta,K,\epsilon_{1},\epsilon_{2}$ (middle left) are adjusted to map the effective Hamiltonian $\hat{H}_{KC}$ to the double-well Hamiltonian modeling the molecular system (middle right). The bottom left shows the Wigner transform phase space representation of the device quantum states, which are analogous to the states in molecular the double-well potential in the bottom right.
Following Venkatraman et al.29 we assume that the noisy dynamics of the device can be described by the following Lindblad equation:
@@
where $\kappa$ sets the photon loss rate corresponding to an effective dissipation rate due to coupling of the system with a surrounding environment and $n_{\text{th}}$ represents the average thermal photon population, a quantity determined by the temperature. Like $\{K,\Delta,\epsilon_{1},\epsilon_{2}\}$, the parameters $\kappa$ and $n_{\text{th}}$ are experimentally tunable and determine the noise associated with the experimental device. 37
@@
which is commonly used to simulate hydrogen bonded complexes. 18 The position and momentum operators of the reaction coordinate, $\hat{x}$ and $\hat{p}$, satisfy the commutation relation $[\hat{x},\hat{p}]=i\hbar$. In the equation above, $m$ represents the effective mass associated with motion along the reaction coordinate. The parameters $\{k_{1},k_{2},k_{4}\}$ are positive and real, typically determined by fitting the ab initio potential energy surface. This fitting process ensures the accurate representation of the barrier height, the curvature of the surface at the reactant and product wells, and the relative stability of reactants and products as parametrized by $k_{1}$. Table 1 lists the parameters $\{k_{1},k_{2},k_{4}\}$ used in this study for numerical simulations of proton transfer dynamics (with $m=1836$ amu) in the four model systems illustrated in Fig. 2.
@@
Adenine-Thymine (DNA) 38 $1.4\times 10^{-3}$ $1.08\times 10^{-2}$ $5.2\times 10^{-3}$
Guanine-Cytosine (DNA) 39 $7.7\times 10^{-4}$ $6.9\times 10^{-3}$ $4.5\times 10^{-3}$
Malonaldehyde (cis-trans) 40, 41 $9.4\times 10^{-5}$ $3.0\times 10^{-3}$ $2.9\times 10^{-3}$
Malonaldehyde (cis-cis) 40, 41 $7.1\times 10^{-4}$ $4.0\times 10^{-3}$ 0
@@
Figure 2: Hydrogen bonded complexes studied to analyze intramolecular proton transfer in cis-cis and cis-trans malonaldehyde (top), intramolecular proton transfer in adenine-thymine (purine and pyrimidine) (middle), and guanine-cytosine (purine and pyrimidine) base pairs.
@@
We now map the parameters $\{\Delta,K,\epsilon_{2},\epsilon_{1}\}$ onto $\{m,k_{1},k_{2},k_{4}\}$ by equating the coefficients of the $\hat{p}^{2}$, $\hat{x}^{4}$, $\hat{x}^{2}$ and $\hat{x}$ terms in Eqs. (3) and (A Roadmap for Simulating Chemical Dynamics on a Parametrically Driven Bosonic Quantum Device). This leads to the following mapping relations:
@@
Clearly, according to Eqs. (6)-(9), the values of $\{\Delta,K,\epsilon_{2},\epsilon_{1}\}$ depend on the value of $c$. We then utilize the flexibility in choosing the value of $c$ to minimize the effect of the additional terms $\hat{p}^{4}$, $\hat{x}^{2}\hat{p}^{2}$ and $\hat{x}^{2}\hat{p}^{2}$ in Eq. (3). As demonstrated in the SI, this requires choosing a value of $c$ small enough to satisfy the following inequality:
@@
Using the cQED device to encode $\hat{H}_{\mathrm{DW}}$ requires a finite, non-zero value for the Kerr non-linearity $K$, so the units of $\Delta$, $\epsilon_{2}$ and $\epsilon_{1}$ are expressed in terms of $K$. For vanishingly small values of $c$, the ratios $\frac{\Delta}{K}$,$\frac{\epsilon_{2}}{K}$, and $\frac{\epsilon_{1}}{K}$ become very large (panels II of Fig. 4), so they are experimentally unfeasible with current cQED platforms. 36 However, there exists a range of values of $c$ for which the parameters are experimentally accessible, while maintaining a useful degree of accuracy for the energies and stationary states. In particular, the Hamiltonian for the cis-cis malonaldehyde is both experimentally accessible and meets the chemical accuracy criterion for stationary states and energies (see Fig. 4A). Chemical systems with asymmetric free energy profiles pose greater challenges due to the higher number of eigenstates required for addressing kinetic questions. Expanding the range of experimental parameters could enable an accurate simulation of these more challenging double-well problems, although it may also risk a breakdown of the effective Hamiltonian approximation.36
@@
Consistency between the energy levels of the $\hat{H}_{KC}$ and $\hat{H}_{DW}$ Hamiltonians is necessary, but not sufficient to ensure that analog simulations on the Kerr-cat device accurately capture the chemical dynamics. This is because the reaction rate is also influenced by the coupling of the reaction coordinate to the thermal bath of nonreactive DOF. Consequently, the actual reaction rate constant depends not only on the eigenvalues and eigenfunctions of the Hamiltonian but also on the temperature and dissipation rate of the bath. Therefore, being able to simulate chemical dynamics on the Kerr-cat device requires that the dynamics of the device be dissipative (noisy), which is indeed the case. The noisy dynamics of Kerr-cat device are described by the Lindblad equation in Eq. (2) with adjustable photon loss rate constant $\kappa$ and temperature set by $n_{th}$.
@@
The complete methodology for the numerical simulation of the open quantum dynamics is outlined in the SI. Here, we clarify that the initial state for dynamics propagation is localized in the reactant well, which is the higher-energy well of the above-mentioned chemical reactions. To generate this state, we diagonalize the system Hamiltonian and select the first eigenstate with more than 50% density in the reactant well. We then apply a sigmoidal filter function to remove the excess density outside of it (see SI for implementation details). This method creates a localized state and avoids the need to identify the critical points of each potential energy surface. By integrating Eq. (A Roadmap for Simulating Chemical Dynamics on a Parametrically Driven Bosonic Quantum Device) with an initial state $\hat{\rho}(0)$ localized in the reactant well, we obtain the time-evolved state $\hat{\rho}(t)$, which can be used to compute the product population at time $t$, as follows:
@@
The reactant-to-product reaction rate constant, $k=1/T_{X}$, is obtained by fitting $P_{P}(t)$ to an exponential decay. The top subpanels of Fig. 5 show the time evolution of $P_{P}(t)$ for the four proton transfer reactions – cis malonaldehyde, cis-trans malonaldehyde, adenine-thymine, and guanine-cytosine – obtained by solving Eq. (A Roadmap for Simulating Chemical Dynamics on a Parametrically Driven Bosonic Quantum Device) for different values of $c$ (solid lines). These top subpanels also include results where the chemical double-well Hamiltonian in Eq. (A Roadmap for Simulating Chemical Dynamics on a Parametrically Driven Bosonic Quantum Device), $\hat{H}_{DW}$, is replaced by the corresponding device Hamiltonian $-\hat{H}_{KC}$ with additional terms $\hat{p}^{4}$, $\hat{x}^{2}\hat{p}^{2}$ and $\hat{p}^{2}\hat{x}^{2}$ (diamonds). The bottom subpanels of Fig. 5 show the dependence of the inverse reaction rate constant, $T_{X}$, on $c$ for both $\hat{H}_{DW}$ (solid line) and $-\hat{H}_{KC}$ (diamonds).
A close examination of the results in Fig. 5 reveals that the dynamics of both the chemical system and the Kerr-cat device align well, with the rate kinetics described by a rate constant. Additionally, the reaction rate constant shows significantly less sensitivity to the value of $c$ compared to the energy levels and eigenfunctions of $\hat{H}_{DW}$ and $\hat{H}_{KC}$. Specifically, the rate constant predicted by the device matches that predicted for the chemical system at values of $c$ as high as 0.4 $\mathrm{a_{0}}$, much higher than the value of $c=0.1\,\mathrm{a_{0}}$ required to match energy levels and eigenfunctions near the top of the barrier. This is likely due to the effect of dissipation, making the overall dynamics less sensitive to small differences in the potential energy surface.
@@
The results in Fig. 5 were obtained with $\kappa=0.1$ and $n_{\text{th}}=0.1$, which are illustrative of the dissipation on the quantum device. However, the observed trends are insensitive to these specific values, as demonstrated by the rate constants obtained for various values of $\kappa$ and $n_{\text{th}}$ shown in Table 2 for $c=0.1\,\mathrm{a_{0}}$. As expected, $T_{X}$ increases (i.e., the reaction slows down) when $\kappa$ and $n_{\text{th}}$ decrease. Nonetheless, the actual values of $T_{X}$ for the chemical double-well and the Kerr-cat device are consistent, regardless of the $\kappa$ and $n_{\text{th}}$ values.
@@
Additional description of the Hamiltonian mapping, the simulation protocol, benchmarks with experimental observations and dynamics figures for different dissipation parameters and computational basis size.
@@ 1 Code and Data Availability
The Python code for the Hamiltonian analysis and dynamical simulations can be found at https://github.com/dcabral00/cQED4ChemDyn. The data files used to generate the figures in the manuscript are also provided at https://github.com/dcabral00/cQED4ChemDyn and hosted at https://doi.org/10.5281/zenodo.13694461.
@@ 1.1 Mapping the Device Hamiltonian to Chemical Double-Well
We consider the effective Hamiltonian for currently experimentally realizable parametrically-driven Kerr-cat cQED devices, which is given by:28, 29, 30, 31, 32, 47
@@
We also take advantage of the aforementioned flexibility in choosing the value of $c$ to minimize the effect of the spurious $\hat{p}^{4}$, $\hat{x}^{2}\hat{p}^{2}$ and $\hat{p}^{2}\hat{x}^{2}$ terms in Eq. (14). As we show below, doing so requires that we choose a value of $c$ small enough so that it satisfies the following inequality:
@@
The fact that the $\hat{p}^{4}$ term scales like $c^{4}$, while the $\hat{p}^{2}$ term scales like $c^{2}$, also implies that the $\hat{p}^{4}$ will become negligible for a sufficiently small value of $c$. In fact, the same inequality, Eq. (22), can be derived by noting that $\frac{1}{2m}\hat{p}^{2}\gg\frac{E_{a}c^{8}}{\hbar^{4}x_{0}^{4}}\hat{p}^{4}$ is equivalent to $\frac{1}{2m}\gg\frac{E_{a}c^{8}}{\hbar^{4}x_{0}^{4}}\hat{p}^{2}$ and that the momentum is maximal when the particle is around the minima, where the potential energy can be approximated as being harmonic. Invoking the virial theorem for the harmonic oscillator, according to which the expectation values of the kinetic energy is equal to that of the potential energy, and noting that $E_{a}$ sets the potential energy scale for the chemical system, we can then estimate $\hat{p}^{2}$ by $2mE_{a}$ in the inequality $\frac{1}{2m}\gg\frac{E_{a}c^{8}}{\hbar^{4}x_{0}^{4}}\hat{p}^{2}$, which turns it into the inequality $\left(\frac{\hbar^{2}x_{0}^{2}}{2mE_{a}c^{4}}\right)^{2}\gg 1$. Thus, satisfying the inequality in Eq. (22), which is equivalent to $\frac{\hbar^{2}x_{0}^{2}}{2mE_{a}c^{4}}\gg 1$, also guarantees that the $\hat{p}^{4}$ term will becomes negligible compared the the $\hat{p}^{2}$ kinetic energy term.
Finally, the same argument would also hold for an asymmetrical double-well since the length and energy scales of the chemical system are not going to be significantly affected by the addition of the asymmetry.